A trader creates a long butterfly spread from put options with strike prices of $160, $170, and $180 per share by trading a total of 20 option contracts (5 contracts at $160, 10 contracts at $170 and 5 contracts at $180). Each contract is written on 100 shares of stock.

The options are worth $22, $28, and $36 per share of stock. What is the value (payoff) of the butterfly spread at maturity as a function of the then stock price?

What is the profit of the butterfly spread at maturity as a function of the then stock price?
Answer #1

1.
=100*(5*MAX(160-S,0)-10*MAX(170-S,0)+5*MAX(180-S,0))

S<160 =100*(5*160-10*170+5*180)=0
S<170 =500S-80000
S<180 =90000-500S
S>180 =0

2.
=100*(5*MAX(160-S,0)-10*MAX(170-S,0)+5*MAX(180-S,0))-5*22*100+10*28*100-5*36*100

S<160 =-5*22*100+10*28*100-5*36*100=-1000
S<170 =500S-80000-5*22*100+10*28*100-5*36*100=501S-801000
S<180 =90000-500S-5*22*100+10*28*100-5*36*100=89900-500S
S>180 =-5*22*100+10*28*100-5*36*100=-1000